If you add three-fourths of a number to the number itself,you get 49. What is the number?
Answer by oberobic(2304)

x = a number
3/4x = 3/4 * x
.
x + 3/4*x = 49
.
multiply both sides by 4 to eliminate the fraction
.
4(x + 3/4*x) = 4*49
.
4x + 3x = 196
.
7x = 196
.
x = 28
.
Done.